# Asynchronous Federated Reinforcement Learning with Policy Gradient Updates: Algorithm Design and Convergence Analysis

### v1

@@ 2. Problem Setup
(1) $V_{\pi}(s)=\mathop{\mathbb{E}}_{a_{t}\sim\pi(\cdot|s_{t}),\atop s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right].$
@@
(2) $Q_{\pi}(s,a)=\mathop{\mathbb{E}}_{a_{t}\sim\pi(\cdot|s_{t}),\atop s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,~{}a_{0}=a\right].$
@@
(3) $\nu_{\pi_{\theta}}(s,a)=(1-\gamma)\mathop{\mathbb{E}}_{s_{0}\sim\rho}\left[\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,~{}a_{t}=a|s_{0},~{}\pi_{\theta})\right],$
@@
(4) $\max_{\theta}J(\theta)\coloneqq\mathop{\mathbb{E}}_{s\sim\rho}\left[V_{\pi_{\theta}}(s)\right].$
The gradient of $J(\theta)$ can be written as (Schulman et al., 2018):
(5) $\nabla_{\theta}J(\theta)=\mathop{\mathbb{E}}_{\tau}\left[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot A_{\pi_{\theta}}(s_{t},a_{t})\right],$
@@
(6) $g(\theta,\tau)=\sum_{t=0}^{\infty}\nabla\log\pi_{\theta}(a_{t}|s_{t})\sum_{h=t}^{\infty}\gamma^{h}r(s_{h},a_{h}).$
Federated policy gradient: We aim to solve the above problem in an FL setting, where $N$ agents collaboratively train a common policy $\pi_{\theta}$. Specifically, each agent collects trajectories and corresponding reward $r(s_{h},a_{h})$ based on its local policy. Then, each agent $i$ estimates $g(\theta_{i},\tau)$ for training the model $\theta$, and the updated models are aggregated at the server. Motivated by the limitations of synchronous model aggregation in terms of time complexity, in the next section, we present our AFedPG methodology that takes an asynchronous approach to solve (4) in a FL setting.
3. Proposed Asynchronous FedPG
Figure 1. An illustration of the asynchronous federated policy gradient updates. Each agent has a local copy of the environment, and agents may collect data according to different local policies. At each iteration, the agent in the yellow color finishes the local process and then communicates with the server, while the other agents keep sampling and computing locally. In the $k$-th global iteration, $\delta_{k}\in\mathbb{N}$ is the delay, $\widetilde{\tau}_{k-\delta_{k}}$ is the sample collected according to the policy $\pi_{\widetilde{\theta}_{k-\delta_{k}}}$, and $d_{k-\delta_{k}}$ is the updating direction calculated from the sample $\widetilde{\tau}_{k-\delta_{k}}$.
@@ 3. Proposed Asynchronous FedPG
In each global iteration of AFedPG, the server applies only one gradient to update the model from the agent who has finished its local computation, while the other $N-1$ agents keep computing local gradients (unapplied gradients) in parallel. In this asynchronous setting, the models used in each agent are outdated, as the server keeps updating the model. Thus, introduce the notion of delay (or staleness) $\delta\in\mathbb{N}$, which measures the difference between the current global iteration and the past global iteration when the agent received the updated model from the server.
Definition 3.2 ().
(Delay) In the $k$-th global iteration, we denote the delay of the applied gradient as $\delta_{k}\in\mathbb{N}$, and the delays of the unapplied gradients as $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}$, where $j_{k}$ denotes the agent that communicates with the server. $\delta_{k}^{i}$ is the difference between the $k$-th iteration and the iteration where the agent $i$ started to compute the latest gradient. In the final $K$-th global iteration, we have $K$ applied gradients $\{\delta_{k}\}_{k=0}^{K-1}$ and unapplied gradients $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}$. The average and maximum delays can be expressed as
(7) $\bar{\delta}~{}\coloneqq~{}\frac{1}{K-1+|\mathcal{C}_{K}|}\big{(}\sum_{k=0}^{K-1}\delta_{k}+\sum_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}\delta_{K}^{i}\big{)},$
and $\delta_{\max}\coloneqq\max(\max_{k}\delta_{k},\max_{i\in\mathcal{C}_{K}}\delta_{K}^{i})$, respectively.
Asynchronous FedPG: Federated learning is a paradigm in which $N$ agents collaboratively train a common global policy with parameters $\theta\in\mathbb{R}^{d}$. As shown in Figure 1, during the training process, agent $i$ collects trajectories in an online behavior, and computes gradients or updating directions using its local trajectories (also known as samples). Then, agent $i$ transmits local gradients to the central server. In particular, in the $k$-th global iteration, AFedPG training consists of the following three steps:
@@
  • Downlink transmission: The server transmits the current global policy parameters $\theta\in\mathbb{R}^{d}$ back to agent $i$ as soon as it finishes the global update.
The server side procedure of AFedPG is shown in Algorithm 1 and the process at the agent side is shown in Algorithm 2. In the $k$-th iteration, the server operates one global update (Steps 4 and 5) as soon as it receives a direction $d_{k-\delta_{k}}$ from an agent with delay $\delta_{k}$. After the global update, the server sends the updated model back to that agent. In Algorithm 2, after receiving the global model from the server, an agent first gets model parameters $\widetilde{\theta}_{k}$ according to Step 2, and then collects samples based on the policy $\pi_{\widetilde{\theta}_{k}}$ (Steps 3). At last, the agent computes the updating direction $d_{k}$, and sends it to the server as soon as it finishes the local process. Overall, all agents conduct local computation in parallel, but the global model is updated in an asynchronous manner as summarized in Figure 1.
Delay-adaptive lookahead and normalized update: In Step 5 of Algorithm 1, to handle the updates with various delays, we use normalized gradients with controllable sizes. Specifically, the error term $\|e_{k}\|$ in Lemma 4.6 is related to $\|\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\|$ and $\|\nabla J(\theta_{k-1})-\nabla J(\theta_{k})\|$. With the smoothness in Lemma 4.4, we are able to bound the error as $\|\theta_{k-1}-\theta_{k}\|=\eta_{k-1}$. The details of the boundaries are shown in (34).
In Step 2 of Algorithm 2 (in blue), we design a delay-adaptive lookahead update, i.e.,
(8) ${\theta}_{k}=(1-\alpha_{k-\delta_{k}}){\theta}_{k-1}+\alpha_{k-\delta_{k}}\widetilde{\theta}_{k}.$
Notably, this is not the momentum method, because the orders are different. The agent collects samples according to $\widetilde{\theta}_{k}$, which only happens in RL. Thus, it is fundamentally different from conventional FL. This mechanism cancels out the second-order Hessian correction terms:
(9) $(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})+\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})\rightarrow 0,$
and thus assists the convergence analysis. The details are shown in Appendix A.2.
@@
0:  MDP $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$; Number of iterations $K$; Step size $\eta_{k}$, $\alpha_{k}$; Initial $\theta_{0}\in\mathbb{R}^{d}$.
@@
4:          Receive $d_{k-\delta_{k}}$ from the agent $i_{k}$.
5:          $\theta_{k+1}\leftarrow\theta_{k}+\eta_{k}\frac{d_{k-\delta_{k}}}{\|d_{k-\delta_{k}}\|}$
@@
7:          Transmit $\theta_{k+1}$ back to the agent $i_{k}$.
8:  end for
8:  ${\theta^{K+1}}$
@@
0:  $\theta_{k}\in\mathbb{R}^{d}$
1:  Receive $\theta_{k}$ from the server.
2:  $\widetilde{\theta}_{k}\leftarrow\theta_{k}+\frac{1-\alpha_{k-\delta_{k}}}{\alpha_{k-\delta_{k}}}(\theta_{k}-\theta_{k-1})$
3:  $\widetilde{\tau}_{k}\sim p(\cdot|\pi_{\widetilde{\theta}_{k}})$
4:  $d_{k}\leftarrow(1-\alpha_{k})d_{k-1}+\alpha_{k}g(\widetilde{\tau}_{k},\widetilde{\theta}_{k})$
5:  When finish computing, transmit ${d}_{k}$ to the server.
5:  ${d}_{k}$
4. Convergence Analysis
In this section, we derive the convergence rate of AFedPG with the following criteria for the global convergence.
Criteria: In this work, we focus on the global convergence, i.e., finding the parameter $\theta$ s.t. $J^{\star}-J(\theta)\leq\epsilon^{\prime}$, where $J^{\star}$ is the optimal expected return.
4.1. Assumptions
In order to derive the global convergence rates, we make the following standard assumptions (Agarwal et al., 2021b; Ding et al., 2020; Liu et al., 2020; Papini et al., 2018; Xu et al., 2020) on policy gradients, and rewards.
Assumption 4.1 ().
  1. (1)
[…]
  2. (2)
    The score function is $M_{h}$-Lipschitz continuous. In other words, $\forall~{}\theta_{i},~{}\theta_{j}\in\mathbb{R}^{d},~{}s\in\mathcal{S},~{}\text{and}~{}a\in\mathcal{A}$, we have
    (10) $\left\|\nabla\log\pi_{\theta_{i}}(a\mid s)-\nabla\log\pi_{\theta_{j}}(a\mid s)\right\|$ $\leq M_{h}\left\|\theta_{i}-\theta_{j}\right\|.$
  3. (3)
These standard assumptions naturally state the reward, first-order and second-order derivatives of the score function are not infinite, and they hold with common practical parametrization methods, e.g., softmax policies. With function approximation $\pi_{\theta}$, the approximation error may not be 00 in practice. Thus, we follow previous works (Liu et al., 2020; Agarwal et al., 2021b; Ding et al., 2022; Huang et al., 2020) with an assumption on the expressivity of the policy parameterization class.
Assumption 4.2 ().
(Function approximation) $\exists~{}\epsilon_{bias}\geq 0$ s.t. for $\forall~{}\theta\in\mathbb{R}^{d}$, the transfer error satisfies
(11) $\mathbb{E}[\big{(}A_{\pi_{\theta}}(s,a)-(1-\gamma)u^{\star}(\theta)^{\top}\nabla\log\pi_{\theta}(a\mid s)\big{)}^{2}]\leq\epsilon_{bias},$
It means that the parameterized policy $\pi_{\theta}$ makes the advantage function $A_{\pi_{\theta}}(s,a)$ approximated by the score function $\nabla\log\pi_{\theta}(a\mid s)$ as the features. This assumption is widely used with Fisher-non-degenerate parameterization. $\epsilon_{bias}$ can be very small with rich neural network parameterization (Wang et al., 2020), and 00 with a soft-max parameterization (Ding et al., 2022).
4.2. Convergence Rates
Based on the above assumptions, the converges rate of AFedPG is given in Theorem 4.3.
@@ 4.2. Convergence Rates
Let Assumption 4.1 and 4.2 hold. With suitable learning rates $\eta_{k}$ and $\alpha_{k}$, after $K$ global iterations, AFedPG satisfy
@@
(12) $J^{\star}-\mathbb{E}[J(\theta_{K})]~{}\leq~{}\mathcal{O}\big{(}{K^{-\frac{2}{5}}}\cdot(1-\gamma)^{-3}\big{)}+\frac{\sqrt{\epsilon_{bias}}}{1-\gamma},$
@@
where $\epsilon_{bias}$ is from (11). Thus, to satisfy $J^{\star}-J(\theta_{K})\leq\epsilon+\frac{\sqrt{\epsilon_{bias}}}{1-\gamma}$, we need $K=\mathcal{O}(\frac{\epsilon^{-2.5}}{(1-\gamma)^{7.5}})$ iterations. As only one trajectory is required in each iteration, the number of trajectories is equal to $K$, i.e., the sample complexity is $\mathcal{O}(\frac{\epsilon^{-2.5}}{(1-\gamma)^{7.5}})$.
@@
Comparison to the synchronous setting: Synchronous FedPG needs $\mathcal{O}(\epsilon^{-2.5})$ trajectories in total, and each agent needs $\mathcal{O}(\frac{\epsilon^{-2.5}}{N})$ trajectories. However, in synchronous FedGP the server has to wait for the slowest agent at each global step, which can still slow down the training process. Let $t_{i}$ denote the time consumption for agent $i=1,\cdots,N$ at local steps with finite values. Then, the waiting time on the server for each step becomes $t_{\max}\coloneqq\max t_{i}$ for FedPG. Our AFedPG approach keeps the same sample complexity as FedPG, but the server processes the global step as soon as it receives an update, speeding up training. Specifically, the average waiting time on the server is $\bar{t}\coloneqq\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}}<\frac{t_{\max}}{N}$ at each step. Thus, the asynchronous FedPG achieves better time complexity than the synchronous approach regardless of the delay pattern. The advantage is significant when $t_{\max}\gg t_{min}$, which occurs in many practical settings with heterogeneous computation powers across different agents. We illustrate the advantage of AFedPG over synchronous FedPG in Figure 2. As the server only operates one simple summation, without loss of generality, the time consumption at the server-side is negligible.
Figure 2. Comparison of time consumptions between synchronous and asynchronous approaches. The circled numbers denote the indices of global steps.
4.3. Key Lemmas
We list some useful lemmas to construct the proof sketch of Theorem 4.3. The details of the proof can be found in Appendix A.
Under Assumption 4.1 on score functions, the following lemma holds based on the results in (Zhang et al., 2020).
Lemma 4.4.
(13) $\|\nabla J(\theta)-\nabla J(\theta^{\prime})\|~{}\leq$ $~{}L_{g}\|\theta-\theta^{\prime}\|,$
Under Assumption 4.1 and 4.2, we utilize the result in (Ding et al., 2022) as the Lemma 4.5.
Lemma 4.5.
(Relaxed weak gradient domination) Under Assumptions 4.1 and 4.2, it holds that
(14) $\|\nabla J(\theta)\|+\epsilon_{g}~{}\geq~{}\sqrt{2\mu}(J^{\star}-J(\theta)),$
where $\epsilon_{g}=\frac{\mu_{F}\sqrt{\epsilon_{bias}}}{M_{g}(1-\gamma)}$ and $\mu=\frac{\mu_{F}^{2}}{2M_{g}^{2}}$.
Based on Lemma 4.4, we derive our milestone, the ascent lemma with delayed updates, as follows:
Lemma 4.6.
(Ascent Lemma with Delay) Under Assumptions 4.1, it holds that
(15) $-J(\theta_{k+1})~{}\leq~{}-J(\theta_{k})-\frac{1}{3}\eta_{k}\|\nabla J(\theta_{k})\|+\frac{8}{3}\eta_{k}\|e_{k}\|+\frac{L_{g}}{2}\eta_{k}^{2},$
where $e_{k}\coloneqq d_{k-\delta_{k}}-\nabla J(\theta_{k})$.
Table 2. Description of the MuJoCo environments.
Tasks Agents Action Space Dimension State Space Dimension
Swimmer-v4 Three-link swimming robot $2$ $8$
Hopper-v4 Two-dimensional one-legged robot $3$ $11$
Humanoid-v4 Three-dimensional bipedal robot $17$ $376$

### v5

@@ 3 Problem Setup
$V_{\pi}(s)=\mathop{\mathbb{E}}_{a_{t}\sim\pi(\cdot|s_{t}),\atop s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right].$ (1)
@@
$Q_{\pi}(s,a)=\mathop{\mathbb{E}}_{a_{t}\sim\pi(\cdot|s_{t}),\atop s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,~{}a_{0}=a\right].$ (2)
@@
$\nu_{\pi_{\theta}}(s,a)=(1-\gamma)\mathop{\mathbb{E}}_{s_{0}\sim\rho}\left[\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,~{}a_{t}=a|s_{0},~{}\pi_{\theta})\right],$ (3)
@@
$\max_{\theta}J(\theta)\coloneqq\mathop{\mathbb{E}}_{s\sim\rho}\left[V_{\pi_{\theta}}(s)\right].$ (4)
The gradient of $J(\theta)$ (Schulman et al., 2018) can be written as:
$\nabla_{\theta}J(\theta)=\mathop{\mathbb{E}}_{\tau}\left[\sum_{t=0}^{\infty}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})\cdot A_{\pi_{\theta}}(s_{t},a_{t})\right],$ (5)
@@
$g(\theta,\tau)=\sum_{t=0}^{\infty}\nabla\log\pi_{\theta}(a_{t}|s_{t})\sum_{h=t}^{\infty}\gamma^{h}r(s_{h},a_{h}).$ (6)
Federated policy gradient: We aim to solve the above problem in an FL setting, where $N$ agents collaboratively train a common policy $\pi_{\theta}$. Specifically, each agent collects trajectories and corresponding reward $r(s_{h},a_{h})$ based on its local policy. Then, each agent $i$ estimates $g(\theta_{i},\tau)$ for training the model $\theta$, and the updated models are aggregated at the server. Motivated by the limitations of synchronous model aggregation in terms of time complexity, in the next section, we present our AFedPG methodology that takes an asynchronous approach to solve equation 4 in an FL setting.
4 Proposed Asynchronous FedPG
Figure 1: An illustration of the asynchronous federated policy gradient updates. Each agent has a local copy of the environment, and agents may collect data according to different local policies. At each iteration, the agent in the yellow color finishes the local process and then communicates with the server, while the other agents keep sampling and computing local gradients in parallel. In the $k$-th global iteration, $\delta_{k}\in\mathbb{N}$ is the delay, $\widetilde{\tau}_{k-\delta_{k}}$ is the sample collected according to the policy $\pi_{\widetilde{\theta}_{k-\delta_{k}}}$, and $d_{k-\delta_{k}}$ is the updating direction calculated from the sample $\widetilde{\tau}_{k-\delta_{k}}$.
@@ 4 Proposed Asynchronous FedPG
In each global iteration of AFedPG, the server applies only one gradient to update the model from the agent who has finished its local computation, while the other $N-1$ agents keep computing local gradients (unapplied gradients) in parallel. In this asynchronous setting, the models used in each agent are outdated, as the server keeps updating the model. Thus, we introduce the notion of delay (or staleness) $\delta\in\mathbb{N}$, which measures the difference between the current global iteration and the past global iteration when the agent received the updated model from the server.
Definition 4.2.
(Delay) In the $k$-th global iteration, we denote the delay of the applied gradient as $\delta_{k}\in\mathbb{N}$ (an integer), and the delays of the unapplied gradients as $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{k}\setminus\{j_{k}\}}$, where $j_{k}$ denotes the agent that communicates with the server. $\delta_{k}^{i}$ is the difference between the $k$-th iteration and the iteration where the agent $i$ started to compute the latest gradient. In the final $K$-th global iteration, we have $K$ applied gradients $\{\delta_{k}\}_{k=0}^{K-1}$ and unapplied gradients $\{\delta_{k}^{i}\}_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}$. The average delay can be expressed as
$\bar{\delta}~{}\coloneqq~{}\frac{1}{K-1+|\mathcal{C}_{K}|}\big{(}\sum_{k=0}^{K-1}\delta_{k}+\sum_{i\in\mathcal{C}_{K}\setminus\{j_{K}\}}\delta_{K}^{i}\big{)}.$ (7)
Asynchronous FedPG: Our goal is to train a global policy with parameters $\theta\in\mathbb{R}^{d}$ via FL across $N$ distributed agents. As shown in Figure 1, during the training process, agent $i$ collects trajectories in an online behavior, and computes gradients or updating directions using its local trajectories (also known as samples). Then, agent $i$ transmits local gradients to the central server. In particular, in the $k$-th global iteration, the training of our AFedPG consists of the following three steps:
@@
  • Downlink transmission: The server transmits the current global policy parameters $\theta\in\mathbb{R}^{d}$ back to the agent $i$ as soon as it finishes the global update.
The server side procedure of AFedPG is shown in Algorithm 1 and the process at the agent side is shown in Algorithm 2. In the $k$-th global iteration, the server operates one global update (Steps 4 and 5) as soon as it receives a direction $d_{k-\delta_{k}}$ from an agent with delay $\delta_{k}$. After the global update, the server sends the updated model back to that agent. In Algorithm 2, after receiving the global model from the server, an agent first gets model parameters $\widetilde{\theta}_{k}$ according to Step 2, and then collects samples based on the policy $\pi_{\widetilde{\theta}_{k}}$ (Steps 3). At last, the agent computes the updating direction $d_{k}$, and sends it to the server as soon as it finishes the local process. Overall, all agents conduct local computation in parallel, but the global model is updated in an asynchronous manner as summarized in Figure 1.
Normalized update at the server: In Step 5 of Algorithm 1, to handle the updates with various delays, we use normalized gradients with controllable sizes. Specifically, the error term $\|e_{k}\|$ in Lemma B.9 is related to $\|\nabla J(\widetilde{\theta}_{k-\delta_{k}})-\nabla J(\widetilde{\theta}_{k})\|$ and $\|\nabla J(\theta_{k-1})-\nabla J(\theta_{k})\|$. With the smoothness in Lemma B.7, we are able to bound the error as $\|\theta_{k-1}-\theta_{k}\|=\eta_{k-1}$. The details of the boundaries are shown in equation 34.
Delay-adaptive lookahead at the agent: In Step 7 of Algorithm 1 (in blue), we design a delay-adaptive lookahead update technique, which is designed specifically for asynchronous FedPG. It operates as follows:
${\theta}_{k}=(1-\alpha_{k-\delta_{k}}){\theta}_{k-1}+\alpha_{k-\delta_{k}}\widetilde{\theta}_{k}.$ (8)
We note that equation 8 is not the conventional momentum method, because the orders are different. Here, $\widetilde{\theta}_{k}$ “looks ahead” based on global policies ${\theta}_{k-1}$ and ${\theta}_{k}$, and the learning rate $\alpha_{k-\delta_{k}}$ depends on the delay $\delta_{k}$.
In equation 8, the agent collects samples according to the current parameter $\widetilde{\theta}_{k}$ in an online behavior, which only happens in the RL setting. This makes the problem fundamentally different from the conventional FL on supervised learning with a fixed dataset. This mechanism cancels out the second-order Hessian correction terms:
$(1-\alpha_{k-\delta_{k}})\nabla^{2}J(\theta_{k})(\theta_{k-1}-\theta_{k})+\alpha_{k-\delta_{k}}\nabla^{2}J(\theta_{k})(\widetilde{\theta}_{k}-\theta_{k})\rightarrow 0,$ (9)
and thus assists the convergence analysis. The details of the derivation are shown in Appendix B.4 marked in blue.
@@
0:  MDP $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$; Number of iterations $K$; Step size $\eta_{k}$, $\alpha_{k}$; Initial $\theta_{0},d_{0}\in\mathbb{R}^{d}$.
@@
4:          Receive $g(\widetilde{\tau}_{k-\delta_{k}},\widetilde{\theta}_{k-\delta_{k}})$ from the agent $i_{k}$.
5:          $d_{k-\delta_{k}}\leftarrow(1-\alpha_{k-\delta_{k}})d_{k-1-\delta_{k-1}}+\alpha_{k-\delta_{k}}g(\widetilde{\tau}_{k-\delta_{k}},\widetilde{\theta}_{k-\delta_{k}})$
@@
7:          $\theta_{k}\leftarrow\theta_{k-1}+\eta_{k-1}\frac{d_{k-\delta_{k}}}{\|d_{k-\delta_{k}}\|}$
8:          $\widetilde{\theta}_{k}\leftarrow\theta_{k}+\frac{1-\alpha_{k-\delta_{k}}}{\alpha_{k-\delta_{k}}}(\theta_{k}-\theta_{k-1})$    # Lookahead
[…]
10:          Transmit $\widetilde{\theta}_{k}$ back to the agent $i_{k}$.
11:  end for
11:  ${\theta^{K}}$
@@
0:  $\widetilde{\theta}_{k^{\prime}}\in\mathbb{R}^{d}$
1:  Receive $\widetilde{\theta}_{k^{\prime}}$ from the server.
2:  $\widetilde{\tau}_{k^{\prime}}\sim p(\cdot|\pi_{\widetilde{\theta}_{k^{\prime}}})$    # Sampling
3:  Estimate policy gradient $g(\widetilde{\tau}_{k^{\prime}},\widetilde{\theta}_{k^{\prime}})$ according to equation 6.
4:  When finish computing, transmit $g(\widetilde{\tau}_{k^{\prime}},\widetilde{\theta}_{k^{\prime}})$ to the server.    # When the server receives the policy gradient, it is the $k$-th step on the server, where $k-{\delta}_{k}=k^{\prime}$.
4:  $g(\widetilde{\tau}_{k^{\prime}},\widetilde{\theta}_{k^{\prime}})$
5 Convergence Analysis
In this section, we derive the convergence rates of AFedPG with the following criterions for the FOSP and global convergence, respectively.
Global Criterion: We focus on the global convergence, i.e., finding the parameter $\theta$ s.t. $J^{\star}-J(\theta)\leq\epsilon^{\prime}$, where $J^{\star}$ is the optimal expected return.
FOSP Criterion: We focus on the first-order stationary convergence, i.e., finding the parameter $\theta$ s.t. $\|\nabla J(\theta)\|\leq\epsilon$.
We use several standard assumptions listed in Appendix B.1. Based on these assumptions, the convergence rates of AFedPG are given in Theorem 5.2 and 5.1.
@@ 5 Convergence Analysis
(Global) Let Assumption B.1 and B.2 hold. With suitable learning rates $\eta_{k}$ and $\alpha_{k}$, after $K$ global iterations, AFedPG satisfies
@@
$J^{\star}-\mathbb{E}[J(\theta_{K})]~{}\leq~{}\mathcal{O}\big{(}{K^{-\frac{2}{5}}}\cdot(1-\gamma)^{-3}\big{)}+\frac{\sqrt{\epsilon_{{\rm bias}}}}{1-\gamma},$ (10)
@@
where $\epsilon_{{\rm bias}}$ is from equation 13. Thus, to satisfy $J^{\star}-J(\theta_{K})\leq\epsilon+\frac{\sqrt{\epsilon_{{\rm bias}}}}{1-\gamma}$, we need $K=\mathcal{O}(\frac{\epsilon^{-2.5}}{(1-\gamma)^{7.5}})$ iterations. As only one trajectory is required in each iteration, the number of trajectories is equal to $K$, i.e., the sample complexity is $\mathcal{O}(\frac{\epsilon^{-2.5}}{(1-\gamma)^{7.5}})$.
Theorem 5.2.
(FOSP) Let Assumption B.1 hold. With suitable learning rates $\eta_{k}$ and $\alpha_{k}$, after $K$ global iterations, AFedPG satisfies
$\mathbb{E}[\|\nabla J(\bar{\theta}_{K})\|]~{}\leq~{}\mathcal{O}\big{(}{K^{-\frac{2}{7}}}\cdot(1-\gamma)^{-3}\big{)},$ (11)
where $\mathbb{E}\|\nabla J(\bar{\theta}_{K})\|\coloneqq\frac{\sum_{k=1}^{K}\eta_{k}\mathbb{E}\|\nabla J(\theta_{k})\|}{\sum_{k=1}^{K}\eta_{k}}$ is the average of gradient expectations. To satisfy $\nabla J(\bar{\theta}_{K})\leq\epsilon$, we need $K=\mathcal{O}(\frac{\epsilon^{-3.5}}{(1-\gamma)^{7.5}})$ iterations. As only one trajectory is required in each iteration, the number of collected trajectories is equal to $K$, i.e., the sample complexity is $\mathcal{O}(\frac{\epsilon^{-3.5}}{(1-\gamma)^{7.5}})$.
@@
Comparison to the synchronous setting: Synchronous FedPG needs $\mathcal{O}(\epsilon^{-2.5})$ trajectories in total, and each agent needs $\mathcal{O}(\frac{\epsilon^{-2.5}}{N})$ trajectories. However, in synchronous FedGP the server has to wait for the slowest agent at each global step, which can still slow down the training process. Let $t_{i}$ denote the time consumption for agent $i=1,\cdots,N$ at local steps with finite values. As the agent has the same computation requirement in each iteration (The number of collected samples is the same.), we assume that the time complexity in each iteration is the same. Then, the waiting time on the server for each step becomes $t_{\max}\coloneqq\max t_{i}$ for FedPG. Our AFedPG approach keeps the same sample complexity as FedPG, but the server processes the global step as soon as it receives an update, speeding up training. Specifically, the average waiting time on the server is $\bar{t}\coloneqq\frac{1}{\sum_{i=1}^{N}\frac{1}{t_{i}}}<\frac{t_{\max}}{N}$ at each step. Thus, the asynchronous FedPG achieves less time complexity than the synchronous approach regardless of the delay pattern. The advantage is significant when $t_{\max}\gg t_{min}$, which occurs in many practical settings with heterogeneous computation powers across different agents. We illustrate the advantage of AFedPG over synchronous FedPG in Figure 5 in Appendix A.1. As the server only operates one simple summation, without loss of generality, the time consumption at the server side is negligible.
6 Experiments
6.1 Setup
Environment: To validate the effectiveness of our approach via experiments, we consider four popular MuJoCo environments for robotic control (Swimmer-v4, Hopper-v4, Walker2D-v4, and Humanoid-v4) (Todorov et al., 2012) with the MIT License. Both the state and action spaces are continuous. Environmental details are described in Table 2 in Appendix A.2, and the MuJoCo tasks are visualized in Figure 2.
Measurement: All convergence performances are measured over $10$ runs with random seeds from 00 to $9$. The solid lines in our main experimental results are the averaged results, and the shadowed areas are confidence intervals with the confidence level $95\%$. The lines are smoothed for better visualization.
Implementation: Policies are parameterized by fully connected multi-layer perceptions (MLPs) with settings listed in Table 3 in Appendix A.1. We follow the practical settings in stable-baselines3 (Raffin et al., 2021) to update models with generalized advantage estimation (GAE) ($0.95$) (Schulman et al., 2018) in our implementation. We use PyTorch (Paszke et al., 2019) to implement deep neural networks (DNNs). All tasks are trained on NVIDIA A100 GPUs with $40$ GB of memory.
Baselines: We first consider the conventional PG approach with $N=1$, to see the effect of using multiple agents for improving sample complexity. We then consider the synchronous FedPG method as a baseline to observe the impact of asynchronous updates on enhancing the time complexity. To see the effect of our delay-adaptive technique, we also consider the performance of AFedPG without the delay-adaptive updates, namely vanilla in Figure 4. Finally, we consider A3C (Mnih et al., 2016), an asynchronous method designed for RL. We note that only a few prior works could be used on the federated PG problem, e.g., A3C, and many existing works in federated supervised learning are not directly applicable to our federated RL setting.
Performance metrics: We consider the following metrics:
  1. 1.
  2. 2.
  3. 3.